# Age-of-Information-Aware Distributed Task Offloading and Resource Allocation in Mobile Edge Computing Networks

### v1

@@ 5.1 Optimality
From Sections 4.1 and 4.2, the RA solutions (22) and (24) are globally optimal for given $\mathbf{X}$. Therefore, there is no optimality loss in the RA solutions. However, in the optimization of $\mathbf{X}$, there may exist a gap between the dual solution and primal solution. Here, we analyze the duality gap of the UA problem and the convergence of the proposed scheme.
@@
$\frac{\partial g}{\partial\mu_{j}}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}^{*}-\frac{\mu_{j}}{2}$ (32)
$\frac{\partial g}{\partial\nu_{j}}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(p),kj}}x_{ij}^{*}-\frac{\nu_{j}}{2}.$
@@
$\frac{\mu_{j}}{2}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}^{*}~{}\text{and}~{}\frac{\nu_{j}}{2}=\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(p),kj}}x_{ij}^{*}.$ (33)
@@
    $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)^{2}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{ij}\right)^{2}$
@@
Let us define the UA variable recovered by the dual solution to Problem $\mathcal{P}_{7}$ as $\hat{\mathbf{X}}$, $\hat{\mathbf{a}}$, and $\hat{\mathbf{b}}$. Then, the performance gap between the solution $\hat{\mathbf{X}}$ and the globally optimal solution is bounded by $\mathcal{O}\left(\left(\Delta_{j}^{(1)}\right)^{2}+\left(\Delta_{j}^{(2)}\right)^{2}\right)$, where $\Delta_{j}^{(1)}=\sqrt{a_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{cm},ij}}x_{kj}\right)$ and $\Delta_{j}^{(2)}=\sqrt{b_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{D_{\text{ES}(s),ij}}x_{ij}\right)$.
@@ 6 Performance Evaluation
In this section, we evaluate the proposed UARA offloading scheme in realistic AI-native MEC systems. Our scheme is tested and compared in diverse network scenarios. Our network scenarios include different network loads, different number of MDs and ESs and different simulation parameters. We first introduce our simulation environment in subsection 6.1, and show the performance of our scheme in subsection 6.2.
@@ 6.1 Simulation Environment
TABLE III: Simulation Settings
@@
Parameters Values Parameters Values $N_{\text{MD}}$ 20 to 160 $N_{\text{ES}}$ 4 to 12 Bandwidth (MHz) 10 Maximum transmission power (dBm) 30 Noise power (dBm/Hz) -174 Pathloss (dB) $41+28log_{10}(d)$ $\delta_{\text{cp}}$ $10^{9}$ $\alpha$ 0 to 100 $\delta_{\text{cm},i}$ 2.6
@@
We use the 3GPP small cell simulation document to model our channel [38, 40]. The parameters are listed in detail in Table III. The cell consists of two type of users: clustered users and uniformly distributed users. Clustered users typically increase the overall latency of the network, as the users in close proximity are more likely to be associated to the same ES. Still, clustered users are considered in order to better mimic the actual network scenarios. These different types of users allow us to simulate a more realistic cell network environment. Additionally, in order to reproduce a more task-intensive network, we prompt our users to generate tasks right after the completion of their previous tasks, which increases the chance of the network saturating.
@@
The number of MDs ranges from 20 to 160. We use various types of MDs, including energy-limited mobile devices and energy-unlimited local computers. For energy-limited mobile devices, we use Samsung’s Galaxy S23, Huawei’s Mate 60, and Apple’s iPhone 14, some of the latest mobile devices available in the market. As for the energy-unlimited computers, we use Apple’s iMac with M1 processor. We use 4 to 12 number of ESs for our simulation. Each ES is equipped with computing resources that outperforms MD’s computing power. We use a combination of NVIDIA’s RTX 2080, RTX 3090 and A6000 as the ES’s computing device.
@@
As for the tasks that the users generate, we use both the language-based AI tasks and computer vision tasks. Typically, large language models (LLMs) that deal with language tasks require large computing power and small input size. In contrast, computer vision models require comparably small computing resources, while the communication loads are relatively big due to larger input data. We combine these two tasks in different ratio. For language tasks, we employ Llama 2 model with approximately 7 billion parameters. As for vision tasks, we use a combination of convolutional neural network (CNN)-based models and ViT-based models. Specifically, we adopt 2 residual net (ResNet) models, 2 MobileNet models, ViT-based side adapter network (SAN) and pyramid scene parsing network (PSPNet). Using these tasks, we simulate three types of networks: network with high communication load, network with heavy computational load and a network with balanced load. In a network with high communication load, the tasks predominantly consist of vision tasks. The task parameters are listed in Table IV. The table values are based on the simulation results and calculations conducted by the authors. As for a network with high computational load, the tasks that the users generate is mainly comprised of language tasks.
TABLE IV: Communication and Computation Loads of the AI Tasks
@@ 6.2.1 Latency Analysis
Figure 4 illustrates the latency of the proposed and the reference schemes with respect to the number of MDs, which ranges from 20 to 160. Our scheme is tested in three network scenarios: network with balanced load, communication-intensive network, and computational-intensive network. The simulation results show that for all network scenarios, the proposed achieves the lowest latency. More specifically, in a network with 160 MDs, the proposed scheme achieves 55.3%, 81.1%, and 64.7% lower latency compared to the best performing reference scheme in each of the network scenario. From the figures, we can see that the proposed scheme achieves a latency that is almost constant compared to the baseline schemes, which grow linearly. From Fig. 4b, which depicts the latency of the schemes in a network with high communication load, we observe that the max SINR scheme outperforms other baselines due to its performance in dealing with communication delay. However, it does not excel the performance of the proposed scheme. Fig. 4c shows the latency of the schemes in a computation-intensive network. Similar to Fig. 4b, max compute scheme achieves the lowest latency among the baseline schemes. Nonetheless, it still does not exceed the performance of the proposed scheme. The reason for the proposed scheme’s exceptional performance is based on the simultaneous consideration of communication/computational latency and flexible offloading policy. How the proposed scheme outperforms other baselines is better illustrated in Figs. 5 and 6.
@@
Figures 5 and 6 show the major latency component of different network scenarios as the number of users increases. Specifically, Fig. 5 illustrates the latency of users in a network with heavy communication load. As for the baseline schemes, the performance degrades severely as the available number of ESs decreases from 8 to 4. However, the proposed scheme maintains similar latency for all scenarios, showing that the algorithm is comparatively invariant to the amount of available resources. In a network with high communication load, the communication latency term becomes the dominating factor in the overall latency of equation (19). Hence, the optimization process of the proposed method is focused on optimizing the communication latency. As a result, the proposed scheme achieves the lowest communication latency. Our method achieves latency lower than the max SINR method, which can be regarded as a theoretical short-term lower bound of the communication latency.
@@ 6.2.2 Energy Analysis
TABLE V: Energy consumption per task of each scheme in different network environment with $N_{\text{ES}}=12$. Local energy consumption is in mWh, and edge energy consumption is in Wh.
Network type Scheme $N_{\text{MD}}$=40 $N_{\text{MD}}$=100 $N_{\text{MD}}$=160 Local Edge Local Edge Local Edge Communication -intensive Random 0.280 24.6 0.231 28.1 0.232 31.0 Max SINR 0.172 19.5 0.128 22.4 0.128 25.2 Max Compute 0.184 27.0 0.260 29.0 0.303 30.8 Combined 0.177 23.6 0.183 26.4 0.211 28.7 Proposed 0.062 8.42 0.061 8.97 0.077 9.44 Computation -intensive Random 0.311 132 0.361 138 0.441 141 Max SINR 0.301 126 0.324 149 0.365 182 Max Compute 0.338 131 0.275 167 0.272 199 Combined 0.315 109 0.303 138 0.332 166 Proposed 0.129 89.9 0.182 94.8 0.241 100
Now, we analyze the energy consumption results of the proposed scheme and the baseline schemes. The energy consumption per task under different network scenarios is shown in Table V. The values indicate that our method achieves the lowest local energy consumption per task for both the communication-intensive network and computation-intensive network. In each network environment, our method consumed up to 64.0% and 57.1% less energy respectively compared to the best performing baseline model. The remarkable energy conservation performance while achieving low latency lies in the strategic offloading policy. As for the communication-intensive networks, since tasks impose high communication load, increasing the local computation ratio proves to be a viable solution. Therefore, the proposed scheme achieves an outstanding performance by modulating its offloading ratio accordingly. Conversely, in computation-intensive networks, where tasks have high computation load with relatively low communication load, increasing the offloading ratio could lead to better performance in both energy conservation and latency. Hence, the proposed method achieves the best performance by adaptively regulating its offloading ratio.
@@ 6.2.3 Pareto Graph Analysis
Figure 7: Pareto efficiency graph of different networks with 80 MDs and 8 ESs, as $\epsilon$ ranges from $0.0$ to $1.0$: (a) Network with high communication load; (c) Network with high computational load.
From Table V, we analyzed how offloading ratio could effect the latency and the energy conservation performance. This leads to a question: Could the baseline schemes outperform the proposed scheme if they increase their local computation ratio? To answer this question, we plot a Pareto efficiency graph of the reciprocals of latency and energy consumption of the models. Pareto efficiency graph illustrates a trade-off relationship between two or more values. We compare the schemes’ performance by examining their locations on the Pareto graph. Since better latency and energy consumption leads to larger reciprocals of their values, if the proposed scheme is located at the upper right of the baseline schemes, this indicates that the proposed scheme excels over the baselines in both latency and energy consumption. Additionally, if the proposed scheme surpasses the baseline in all axes, it indicates that the proposed scheme Pareto-dominates the baseline, signifying its superiority in all aspects.
@@ 6.2.4 Effects of unit-changing parameter $\alpha$
Figure 8: Average latency and energy consumption of the network for various $\alpha$ ($N_{\text{ES}}=4$): LABEL:sub@subfig:mu_latency Average latency for different $\alpha$ values. Latency rises with an increase of $\alpha$ due to the escalation of local computation; (b) Local energy consumption for different $\alpha$ values. Energy consumption decreases as $\alpha$ rises, due to the increase of local computation.

### v2

@@ 4.1 Optimality
From Sections 3.1 and 3.2, the RA solutions (23) and (25) are globally optimal for given $\mathbf{X}$. Therefore, there is no optimality loss in the RA solutions. However, in the optimization of $\mathbf{X}$, there may exist a gap between the dual solution and primal solution. Here, we analyze the duality gap of the UA problem and the convergence of the proposed scheme.
@@
$\frac{\partial g}{\partial\mu_{j}}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}^{*}-\frac{\mu_{j}}{2}$ (33)
$\frac{\partial g}{\partial\nu_{j}}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(p),kj}}x_{ij}^{*}-\frac{\nu_{j}}{2}.$
@@
$\frac{\mu_{j}}{2}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}^{*}~{}\text{and}~{}\frac{\nu_{j}}{2}=\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(p),kj}}x_{ij}^{*}.$ (34)
@@
    $\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)^{2}+\sum_{j\in\mathcal{E}}\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{ij}\right)^{2}$
@@
Let us define the UA variable recovered by the dual solution to Problem $\mathcal{P}_{7}$ as $\hat{\mathbf{X}}$, $\hat{\mathbf{a}}$, and $\hat{\mathbf{b}}$. Then, the performance gap between the solution $\hat{\mathbf{X}}$ and the globally optimal solution is bounded by $\mathcal{O}\left(\left(\Delta_{j}^{(1)}\right)^{2}+\left(\Delta_{j}^{(2)}\right)^{2}\right)$, where $\Delta_{j}^{(1)}=\sqrt{a_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{cm},ij}}x_{kj}\right)$ and $\Delta_{j}^{(2)}=\sqrt{b_{j}}-\left(\sum_{i\in\mathcal{M}}\sqrt{\hat{D}_{\text{ES}(s),ij}}x_{ij}\right)$.
@@ 4.2 Convergence Analysis
4.3 Computational Complexity
As previously mentioned, the proposed algorithm requires $\mathcal{O}(1/\epsilon^{2})$ steps to converge, with the total computational complexity of each step being $\mathcal{O}(N_{\text{ES}}N_{\text{MD}})$. Therefore, for optimization with a fixed $\epsilon$, the proposed algorithm necessitates total $\mathcal{O}(\frac{1}{\epsilon^{2}}N_{\text{ES}}N_{\text{MD}})$ floating-point operations. By virtue of the distributed optimization, each of MDs and ESs has computational complexities of $O(N_{\text{ES}})$ and $O(N_{\text{MD}})$ in each step, respectively. In contrast, the heuristic exhaustive search method has a complexity of $\mathcal{O}(N_{\text{ES}}^{N_{\text{MD}+1}}N_{\text{MD}})$. In practical scenarios, the number of MDs typically exceeds 20, significantly increasing the computational complexity. Thus, the proposed algorithm effectively reduces computational complexity, making it more efficient and scalable. This efficiency makes it well-suited for practical applications in large-scale MEC systems.
@@ 5 Performance Evaluation
In this section, we evaluate the proposed UARA offloading scheme in realistic AI-native MEC systems. Our scheme is tested and compared in diverse network scenarios. Our network scenarios include different network loads, different number of MDs and ESs and different simulation parameters. We first introduce our simulation environment in subsection 5.1, and show the performance of our scheme in subsection 5.2. For reproducibility of the results, our implementation code is available at https://github.com/mwkim00/mec_uara.
@@ 5.1 Simulation Environment
TABLE II: Simulation Settings
@@
Parameters Values Number of MDs, $N_{\text{MD}}$ 20 to 160 Number of ESs, $N_{\text{ES}}$ 4 to 20 Bandwidth (MHz) 10 Maximum transmission power (dBm) 30 Noise power (dBm/Hz) -174 Pathloss (dB) $41+28\log_{10}(d)$ Coefficient of energy consumption for computing, $\delta_{\text{cp}}$ $10^{9}$ Coefficient of energy consumption for communication, $\delta_{\text{cm},i}$ 2.6 Weight for energy penalized term, $\alpha$ 0 to 100
@@
We use the 3GPP small cell simulation document to model our channel [16, 20]. The parameters are listed in detail in Table II. The cell consists of two type of users: clustered users and uniformly distributed users. Clustered users typically increase the overall latency of the network, as the users in close proximity are more likely to be associated to the same ES. Still, clustered users are considered in order to better mimic the actual network scenarios. These different types of users allow us to simulate a more realistic cell network environment. Additionally, in order to reproduce a more task-intensive network, we prompt our users to generate tasks right after the completion of their previous tasks, which increases the chance of the network saturating.
@@
The number of MDs ranges from 20 to 160. We use various types of MDs, including energy-limited mobile devices and energy-unlimited local computers. For energy-limited mobile devices, we use Samsung’s Galaxy S23, Huawei’s Mate 60, and Apple’s iPhone 14, some of the latest mobile devices available in the market. As for the energy-unlimited computers, we use Apple’s iMac with M1 processor. We use 4 to 20 number of ESs for our simulation. Each ES is equipped with computing resources that outperforms MD’s computing power. We use a combination of NVIDIA’s RTX 2080, RTX 3090 and A6000 as the ES’s computing device.
@@
As for the tasks that the users generate, we use both the language-based AI tasks and computer vision tasks. Typically, large language models (LLMs) that deal with language tasks require large computing power and small input size. In contrast, computer vision models require comparably small computing resources, while the communication loads are relatively big due to larger input data. We combine these two tasks in different ratio. For language tasks, we employ Llama 2 model with approximately 7 billion parameters. As for vision tasks, we use a combination of convolutional neural network (CNN)-based models and ViT-based models. Specifically, we adopt 2 residual net (ResNet) models, 2 MobileNet models, ViT-based side adapter network (SAN) and pyramid scene parsing network (PSPNet). Using these tasks, we simulate three types of networks: network with high communication load, network with heavy computational load and a network with balanced load. In a network with high communication load, the tasks predominantly consist of vision tasks. The task parameters are listed in Table III. The table values are based on the simulation results and calculations conducted by the authors. As for a network with high computational load, the tasks that the users generate is mainly comprised of language tasks.
TABLE III: Communication and Computation Loads of the AI Tasks
@@ 5.2.1 Latency Analysis
A network employing distributed optimization, including the proposed method, is expected to show robustness with regard to the environment as scalability is crucial in distributed optimization. We test each scheme in different network environments to investigate each scheme’s scalability. Figure 5 illustrates the latency of the proposed and the reference schemes with respect to the number of MDs, which ranges from 20 to 160. Our scheme is tested in three network scenarios: network with balanced load, communication-intensive network, and computational-intensive network. The simulation results show that for all network scenarios, the proposed achieves the lowest latency. More specifically, in a network with 160 MDs, the proposed scheme achieves 55.3%, 81.1%, and 64.7% lower latency compared to the best performing reference scheme in each of the network scenario. From the figures, we can see that the proposed scheme achieves a latency that is almost constant compared to the baseline schemes, which grow linearly. From Fig. 5b, which depicts the latency of the schemes in a network with high communication load, we observe that the max SINR scheme outperforms other baselines due to its performance in dealing with communication delay. However, it does not excel the performance of the proposed scheme. Fig. 5c shows the latency of the schemes in a computation-intensive network. Similar to Fig. 5b, max compute scheme achieves the lowest latency among the baseline schemes. Nonetheless, it still does not exceed the performance of the proposed scheme. The reason for the proposed scheme’s exceptional performance is based on the simultaneous consideration of communication/computational latency and flexible offloading policy. How the proposed scheme outperforms other baselines is better illustrated in Figs. 6 and 7.
@@
Figures 6 and 7 show the major latency component of different network scenarios as the number of users increases. Specifically, Fig. 6 illustrates the latency of users in a network with heavy communication load. As for the baseline schemes, the performance degrades severely as the available number of ESs decreases from 8 to 4. However, the proposed scheme maintains similar latency for all scenarios, showing that the algorithm is comparatively invariant to the amount of available resources. In a network with high communication load, the communication latency term becomes the dominating factor in the overall latency of equation (19). Hence, the optimization process of the proposed method is focused on optimizing the communication latency. As a result, the proposed scheme achieves the lowest communication latency. Our method achieves latency lower than the max SINR method, which can be regarded as a theoretical short-term lower bound of the communication latency.
@@
Figure 8: The system latency of networks where $N_{\text{MD}}=60$ and $\alpha=1$: (a) Network with high communication load; (b) Network with high computational load.
Previously, we have observed how different method performs as the number of MDs increases. Next, we investigate the scalability of the method concerning the number of ESs. Figure 8 exhibits the latency of each method relative to the number of ESs. Starting from $N_{\text{ES}}=4$, we gradually increase the number of ESs up to 20 to analyze how the proposed method scales. The results indicate that as the number of ESs increases, the latency of all the methods gradually decreases. More importantly, although necessities of optimization is decreased as more ESs can reduce load their individual load on computing/computation, the proposed method consistently outperforms the baseline methods. This demonstrates that the proposed method is robust to the change of the number of ESs, showing good scalability property while achieving the lowest latency among the schemes.
@@ 5.2.2 Energy Analysis
TABLE IV: Energy consumption per task of each scheme in different network environment with $N_{\text{ES}}=12$. Local energy consumption is in mWh, and edge energy consumption is in Wh.
Network type Scheme $N_{\text{MD}}$=40 $N_{\text{MD}}$=100 $N_{\text{MD}}$=160 Local Edge Local Edge Local Edge Communication -intensive Random 0.280 24.6 0.231 28.1 0.232 31.0 Max SINR 0.172 19.5 0.128 22.4 0.128 25.2 Max Compute 0.184 27.0 0.260 29.0 0.303 30.8 Combined 0.177 23.6 0.183 26.4 0.211 28.7 Proposed 0.062 8.42 0.061 8.97 0.077 9.44 Computation -intensive Random 0.311 132 0.361 138 0.441 141 Max SINR 0.301 126 0.324 149 0.365 182 Max Compute 0.338 131 0.275 167 0.272 199 Combined 0.315 109 0.303 138 0.332 166 Proposed 0.129 89.9 0.182 94.8 0.241 100 * bold & underline: best performance
Now, we analyze the energy consumption results of the proposed scheme and the baseline schemes. The energy consumption per task under different network scenarios is shown in Table IV. The values indicate that our method achieves the lowest local energy consumption per task for both the communication-intensive network and computation-intensive network. In each network environment, our method consumed up to 64.0% and 57.1% less energy respectively compared to the best performing baseline model. The remarkable energy conservation performance while achieving low latency lies in the strategic offloading policy. As for the communication-intensive networks, since tasks impose high communication load, increasing the local computation ratio proves to be a viable solution. Therefore, the proposed scheme achieves an outstanding performance by modulating its offloading ratio accordingly. Conversely, in computation-intensive networks, where tasks have high computation load with relatively low communication load, increasing the offloading ratio could lead to better performance in both energy conservation and latency. Hence, the proposed method achieves the best performance by adaptively regulating its offloading ratio.
@@ 5.2.3 Pareto Graph Analysis
Figure 9: Pareto efficiency graphs of different networks with 80 MDs and 8 ESs, as $\epsilon$ ranges from $0.0$ to $1.0$: (a) Network with high communication load; (b) Network with high computational load.
From Table IV, we analyzed how offloading ratio could effect the latency and the energy conservation performance. This leads to a question: Could the baseline schemes outperform the proposed scheme if they increase their local computation ratio? To answer this question, we plot a Pareto efficiency graph of the reciprocals of latency and energy consumption of the models. Pareto efficiency graph illustrates a trade-off relationship between two or more values. We compare the schemes’ performance by examining their locations on the Pareto graph. Since better latency and energy consumption leads to larger reciprocals of their values, if the proposed scheme is located at the upper right of the baseline schemes, this indicates that the proposed scheme excels over the baselines in both latency and energy consumption. Additionally, if the proposed scheme surpasses the baseline in all axes, it indicates that the proposed scheme Pareto-dominates the baseline, signifying its superiority in all aspects.
@@ 5.2.4 Effects of unit-changing parameter $\alpha$
Figure 10: Average latency and energy consumptions of the network for various $\alpha$ ($N_{\text{ES}}=4$): LABEL:sub@subfig:mu_latency Average latency for different $\alpha$ values. Latency rises with an increase of $\alpha$ due to the escalation of local computation; (b) Local energy consumption for different $\alpha$ values. Energy consumption decreases as $\alpha$ rises, due to the increase of local computation.
@@
5.2.5 Effects of step size $\eta$ on convergence
In Section 4.2, we have analyzed how the sub-gradient descent update step size $\eta$ could affect the convergence of the solution. More specifically, our analysis suggests that by choosing the optimal step size, the optimality gap will monotonically decrease as the iteration $T$ increases. Here, we experimentally visualize how different step sizes affect the update process of Lagrangian multipliers $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$. Finally, we validate that different step sizes could lead to different convergence properties of the proposed scheme.
Figure 11: Convergence graph of the Lagrangian multipliers $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ at different step sizes, where the network consists of 80 MDs and 4 ESs: (a) Lagrange multiplier $\mu_{1}$ for ES 1; (b) Lagrange multiplier $\nu_{1}$ for ES 1.
Figure 12: Convergence graph of the objective functions at different step sizes, where the network consists of 80 MDs and 4 ESs: (a) Primal and dual function values of ES 1; (b) Duality gap of ES 1.
Figure 11 illustrates the effect of different step sizes on the convergence of the Lagrangian multipliers $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$. Although the experimented step sizes all lead to the gradual convergence of the Lagrangian multipliers, some step sizes showed better convergence performance compared to other values. As for Fig. 11, $\eta=0.01$ and $\eta=0.05$ showed fast convergence compared to $\eta=0.001$. Figure 12 exhibits the value of the primal and dual functions of the first ES, as the Lagrangian multipliers $\mu_{1}$ and $\nu_{1}$ updates. Similar to the simulation settings of Fig. 11, we compare the results using step sizes of $0.001$. $0.01$ and $0.05$. These step sizes are applied to both $\eta_{1}$ and $\eta_{2}$. Consistent with the previous results of Fig. 11, $\eta=0.01$ and $\eta=0.05$ demonstrated a good convergence speed. From Fig. 12a, it can be observed that the primal and dual functions corresponding to $\eta=0.01$ has the smallest duality gap. It can be inferred that a step size of $0.01$ approximates the optimum more closely, whereas a step size of $0.05$ fails to achieve a similar level of precision. The step size of $0.001$ showed the slowest convergence, due to its small step size. The duality gap of 10 implies that one MD experienced approximately $10/80=0.125$ seconds of delay difference compared to the ideal latency. The difference is negligible, given that the average delay of the network was approximately 10 seconds, as can be observed from Fig. 5a. Therefore, we conclude that different step sizes affect the rate at which the solution reaches its optimum.
6 Related Works
6.1 Conventional Offloading Approaches: Pricing-Based Approach
Several studies [19, 24, 25, 26, 27, 28] have been proposed for service/load offloading in cellular networks, characterized by ultra-dense networks, massive MIMO networks, heterogeneous networks (HetNets), and cloud random access networks (C-RAN). In particular, in [1], the authors presented an optimization-based approach to minimize delay in multi-cell MEC systems. Assuming divisible tasks, the solution allocated portions to each ES to minimize delay while considering an energy constraint. In [29], the authors proposed a modern approach for network load balancing. The aim of the approach is distributed optimization of UARA in HetNets. In the work, the concept of pricing-based load balancing has been proposed for the first time, which enables distributed network optimization without any additional information exchange between ESs. However, such approach was limited to an objective function, specifically proportional fairness, where an artistic formulation led to a pricing framework. For instance, in [30], a pricing-based UARA optimization is further extended to backhaul-limited cellular networks.
6.2 Task Offloading in MEC
6.3 Age-of-Information (AoI) minimization
The concept of age-of-information (AoI) is a quantitative metric for the freshness of data, defined as time elapsed since the data was generated. In Remark 3, we demonstrate that the objective function in our proposed system model is equivalent to AoI minimization. Because AoI is closely related to latency minimization, we here introduce the related studies regarding AoI.
In [46], a resource optimization problem for energy harvesting networks is addressed using a difference of convex method to minimize AoI. The study in [47] tackles a joint user association and resource allocation (UARA) problem by transforming a discrete user association variable into a stochastic policy, aiming to maximize the sum-rate under an AoI constraint. Similarly, the authors of [48] propose an AoI-based offloading method for single ES networks, where a stochastic policy is used to minimize AoI. Further, the study in [49] optimizes single-ES task offloading for opportunistic channel access. Multi-base station (BS) task offloading is explored in [50] using a deep Q-network (DQN) approach in a centralized setting. In [51], a multi-BS task offloading method is proposed, where a location-based pricing vector is used to minimize monetary cost under an AoI constraint, employing simulated annealing without resource optimization.
TABLE V: Interpretation of the Notations Used in this Paper